OpenStudy (anonymous):

Help! Kelly's goal is to average 25,00 per month in sales for the first three months of the year. Her sales for January are 40,000 and February are 20,000..What do I do first?January 40,000 February 20,000 and march??

OpenStudy (anonymous):

you will need to find what 3 numbers average to 25,000 assuming two of the numebrs are 20,000 and 40,000. to find average, you do: (20,000 + 40,000 + x)/3 = 25,000 x is the next month. can you solve for x from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you need to multiply both sides by 3 to get rid of the denominator, so now you have: 60,000 + x = 75,000 can you solve from there?

OpenStudy (anonymous):

answer is 15,00 ..Don't know how to work problem? show me

OpenStudy (anonymous):

what I did to solve that was first to think about how I would figure out average. Do you know what that is? watever, if you don't, its adding all the numbers you want to get the average of, then dividing by how many numbers there are. I took 20,000, 40,000 and x(the number we dont know) and found the average, which was supposed to be 25,000 (the goal). then i solved for x. \[\frac{ 20,000+40,000+x }{ 3 }\]and set equal to 25,000

OpenStudy (anonymous):

Thanks..Lost in math..You a life saver

OpenStudy (anonymous):

lol no problem, math can be annoying

OpenStudy (anonymous):

Thanks and God bless you:)